(please show work)

Chapter 02, Problem 08 GO XIncorrect. An 18-year-old runner can complete a 10.0-km course with an average speed of 4.46 m/s. A 50-year-old runner can cover the same distance with an average speed of 4.10 m/s. How much later (in seconds) should the younger runner start in order to finish the course at the same time as the older runner? Numbe Unit No units the tolerance is +/-290 Click if you would like to Show Worl stion: Open Show Work m/s m/s 2 cm/s 2 LINK TO TEXT LINK TO TEXT
Answer #1

t_y=\frac {d}{v_y} (Time taken by the younger guy)

t_o=t_y+t=\frac {d}{v_o} (Time taken by the older guy must be equal to the time taken by the younger guy)

\frac {d}{v_y}+t=\frac {d}{v_o}

t=\frac {d}{v_o}-\frac {d}{v_y}=10000\left (\frac {1}{4.10}-\frac {1}{4.46} \right )\: s= 196.87\: s (Required answer)